OpenStudy (anonymous):

find inverse laplace transform:

OpenStudy (anonymous):

\[\frac{ (s-1)e^{-\pi s/2}}{ (s-1)^2+4 }\]

OpenStudy (loser66):

it forms the form of \(\Large e^{(-\pi/2) *s}F(s)\) where F(s) = \(\dfrac{s-1}{(s-1)^2+4})\ And we get it from the formula #13 on the table

OpenStudy (primeralph):